# The human-machine paradox:

### v5

@@ 4 Design of experiment
In our Monte Carlo experimental setting, we focus on a single firm ($N=1$) walking through ten epochs ($E=10$). With regards to the number of replications, which remarkably impacts the computational effort required by simulations, we initially set $K=1000$, checking the margin of error at 95% confidence level before evaluating an eventual increase of $K$. Then, for the function $a$, we implement three options:
@@
where $\theta_{HM}$ is a notational simplification for $\theta_{i,c=HM,d,e},\forall(i,d,e)$. Herein, meanmean represents the case of simply averaging the abilities of human and machine skills with no augmenting effects, while collaboratecollaborate and superpowersuperpower represent moderate and strong augmentation, respectively. For the function $g^{\prime}$, we deploy two options:
@@
where $y$ is a notational simplification for $y_{i,c,d,e}$ and $\theta$ is a notational simplification for $\theta_{i,c,d,e}$. These options reflect two contrasting dynamics, offering performance-to-output maps. Then, we set $\gamma_{HM}\in[1,2]$, thus limiting the augmenting factor to increase only, in the range [0%,100%]. Also, we impose $mc_{H}\in[0.4,0.6]$ and $mc_{M}\in[0.7,0.9]$, thus putting us in the case of automation bringing higher margin of contribution than human labor, and $mc_{HM}=1-(1-mc_{H})-(1-mc_{M})$ to represent the idea that both labor and capital allocations are required when deploying human-machine collaboration. Finally, we set $\delta_{H}=0$ and $\delta_{M}=0$ as static default, representing the case of human and machine skills being already on their plateau of efficiency. According to these initial constraints, we formulate our experimental design for numeric variables by combining the value selection optimized through the Latin Hypercube Sampling method (LHS) (McKay et al., 2000) and the Maximin method (Johnson et al., 1990) -see Figure 3 for an overview of the resulting pairwise value selection. With regards to the distributions of skill policies and task difficulties, we set $P_{c}\sim U\{P_{H},P_{HM},P_{M}\}$ and $P_{d}\sim U\{P_{Low},P_{Med},P_{High}\}$, i.e. uniformly distributed, and we opt for $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$ for the probability distributions of performance. More in depth, in a first experiment we set the parameter pairs $(\alpha,\beta)_{c,d,e=0}$ and $(\alpha,\beta)_{c,d,e=E-1}$ configuring improvement for human skills across all task difficulties, with static machine skills, as illustrated in Figure 4(a). In a second experiment, which is meant to stress test the simulation results under the effect of improved machine skills, we set the parameter pairs $(\alpha,\beta)_{c,d,e=0}$ and $(\alpha,\beta)_{c,d,e=E-1}$ as illustrated in Figure 4(b). All the experiments are executed on computing instances deploying Python 3.11, generating a total of 57.6 million task executions and 288 million data logs for $\lambda_{i,c,d,e,k}$ with $\lambda\in\{\theta,y,v,err,u\}$.
@@
Figure 4: Distributions of probabilities $P(\theta)_{c,d,e}$ for all skill policies and difficulty levels at epoch $e=0$ and e=E1e=E-1, representing human skills to evolve through epochs, while machine skills are static
@@ 5 Results
Contrary to the conventional wisdom that a human-in-the-loop policy is a safe risk-mitigating strategy, our simulations reveal it prone to contextual economic volatility. As shown in Figure 5, in tasks requiring high generalization capabilities, the human-machine skill policy over performs both the human-exclusive and the machine-exclusive strategies in terms of bottom-line utility, but only when achieving augmenting effects (central tendency of mean utility up to 0.22 for the human-machine team vs 0.08 and 0.03 for the human and machine alone, respectively). Instead, when failing to produce augmenting effects, the human-machine strategy often results the least economically effective, sometimes entering the dangerous zone of negative utility, destroying value instead of building it. Interestingly, the advantage of the human-machine skill policy achieving augmenting effects in high-difficulty tasks is confirmed also when machine skills surpass the performance of human skills across all task difficulties if the context is characterized by high cost of errors (experiment 2, with $c_{err}\geq 0.80$ and $t_{err}\leq 0.10$). While denoting the emergent critical finding regarding the all-or-nothing nature of the human-machine skill policy, our results successfully replicate two relevant established stylized facts from the existing literature. First, the simulations show that automation tends to be the dominant economic strategy when task difficulty is low-to-medium, underscoring its superiority for repetitive or compartmentalized tasks. Second, the framework confirms that the human-exclusive skill policy can be highly competitive, and often superior, when surprises at inference time are the norm and generalization difficulty is high.
(a) Performance ($\Lambda=\Theta$)
(b) Utility ($\Lambda=U$)
Figure 5: Mean of the distributions $\mu(\Lambda^{s}_{c,d})$ with $\Lambda\in\{\Theta,U\}$, with breakdown by experiment and by human-machine interaction dynamics (function $a$)
Moving from overview to a more fine-grained detail, we consolidate three characterizations of our results, with distinction between Experiment 1 and 2, including the interesting close-up on the data of Experiment 2 with focus on the conditions characterized by high cost of error. First, in Table 1 we present a synthetic view over the central tendency for the distributions $\omega(\Lambda^{s}_{[c],[d]})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ under the assumption of uniformly-distributed probability of occurrence for configuration setups, with incremental breakdown by skill policy and task difficulty. Then, in Table 2, we extend the statistical description of these distributions by reporting their standard deviation, range, and median value; additionally, we provide the breakdown by human-machine interaction dynamics (function $a$) for the human-machine skill policy. Next, we visualize the distributions $\omega(\Lambda^{s}_{[c],[d]})$ with $\omega\in\{\mu,\sigma\}$ and $\Lambda\in\{\Theta,Y,V,ERR,U\}$ based on overlapping violin and box plots, (see Figures 6 and 7).
Table 1: Central tendency (arithmetic mean) of the distributions $\omega(\Lambda^{s}_{[c],[d]})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$
Table 2: Statistical description (mean, standard deviation, range, and median) of the distributions $\omega(\Lambda_{[c],[d]}^{s})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$
(a) Exp. 1
(b) Exp. 2
(c) Exp. 2*
Figure 6: Visualization based on overlapping violin and box plots of the distributions $\mu(\Lambda_{[c],[d]}^{s})$ with $\Lambda\in\{\Theta,Y,V,ERR,U\}$ by experiment
(a) Exp. 1
(b) Exp. 2
(c) Exp. 2*
Figure 7: Visualization based on overlapping violin and box plots of the distributions $\sigma(\Lambda_{[c],[d]}^{s})$ with $\Lambda\in\{\Theta,Y,V,ERR,U\}$ by experiment
@@ 6 Discussion
The visualization of the distributions $\mu(\Theta^{s}_{[c],[d]})$ and $\mu(U^{s}_{[c],[d]})$ resulting from both our experiments (see Figure 8) complements the overall effect revealed by the analysis of central tendencies. When the machine skills achieve remarkable performance across all task difficulties (Exp. 2), a straightforward effect is an overall dominance of the machine-exclusive skill policy in terms of utility, with consequential benefits also for the human-machine skill policy -see Figure 8(e). However, as detailed in Figure 8(f), which focuses the lens on situations characterized by potentially high cost of error, when task difficulty is high the human-machine skill policy achieving augmentation still prevails in terms of utility over the machine-exclusive policy. This reinforces the idea that the understanding of the context is crucial for optimal skill policy decisions. Moreover, when narrowing the analysis to contrasting scenarios of cost effectiveness, we observe that an increase over epochs of the margin of contribution associated to the human-machine skill policy mitigates the erosion of value, while not safely resolving it (see Figure 9, with the human-machine skill policy failing to achieve augmentation closing just at break-even). This highlights the importance of the augmenting effect towards the bottom-line economic effectiveness of the human-machine skill policy, even when machine skills do a great job in high-difficulty tasks. With regards to dispersion, the mean of the distributions $\sigma(U^{s}_{[c],[d]})$ and $\rho(U^{s}_{[c],[d]})$ presented in Figure 10 highlights a lower variability for the values in $U^{s}_{[c],[d]}$ over the simulation space. In contrast, in Figure 11 we observe that the standard deviation and the range of the distribution $\mu(U^{s}_{[c],[d]})$ for the human-machine skill policy tends to be slightly higher than than the one realized by the human or machine skills alone. Herein, on both the observations, we note that augmenting effects seem not to have a visible impact. Then, our interpretation is that the hybrid skill strategy contributes more to reducing the variability of utility over the spectrum of scenarios under consideration, while slightly but less evidently increasing the variability of $\mu(U^{s}_{[c],[d]})$. This may explain why the combination of human and machine skills can be perceived as a risk-mitigating strategy (the reduction of variability naturally suggests lower risks); however, the variability alone is not sufficient to avoid the risk of being near or below utility break even. To allow the reader a more precise understanding of the margins of error of $\mu(\Theta^{s}_{[c],[d]})$ and $\mu(U^{s}_{[c],[d]})$ obtained with $K=1000$ as estimates of the true mean ($K\rightarrow\infty$), for each simulation $s$ we compute the estimated MOE at 95% confidence level for $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$ and $\mu(\Lambda^{s}_{c,d})$ with $\Lambda\in\{\Theta,U\}$. Formally, $\forall s$, we compute $\text{MOE}_{95}\big(\mu(\Lambda^{s})\big)=1.96*\sigma(\Lambda^{s})/\sqrt{K^{s}}$, $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c})\big)=1.96*\sigma(\Lambda^{s}_{c})\sqrt{K^{s}_{c}}$, and $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c,d})\big)=1.96*\sigma(\Lambda^{s}_{c,d})\sqrt{K^{s}_{c,d}}$ with $K^{s}=K*E=10000$, and $K^{s}_{c}$ and $K^{s}_{c,d}$ globally and prudentially set to $3000<K*E/|C|\sim 3333$ and $1000<K*E/(|C|*|D|)\sim 1111$, respectively. We acknowledge that, within each simulation $s$, the sets $\Lambda^{s}$, $\Lambda^{s}_{c}$ and $\Lambda^{s}_{c,d}$ pool values that are related to different epochs $e$, thus potentially in relationship with different Beta distributions for skill policies $H$ and $HM$ (this is due to the deterministic interpolation of the parameters $\alpha$ and $\beta$ over epochs for the skill $H$), meaning that the data is heterogeneous and consists of independent samples of smaller size. However, the margins of error computed on the pooled standard deviation represent a conservative upper bound for the granular error that should be selectively computed for each epoch, because the pooled standard deviation combines the actual sampling uncertainty with the additional variability introduced by the parameter interpolation. The distributions of $\text{MOE}_{95}\big(\mu(\Lambda^{s})\big)$, $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c})\big)$ and $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c,d})\big)$ are described in Figure 13 and in Table 13, while Figure 13 presents the the means $\mu(\Lambda^{s}_{[c],[d]})$ against the related $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{[c],[d]})\big)$.
(a) $\mu(\Theta^{s}_{[c],[d]})$, Exp. 1
(b) $\mu(\Theta^{s}_{[c],[d]})$, Exp. 2
(c) $\mu(\Theta^{s}_{[c],[d]})$, Exp. 2*
(d) $\mu(U^{s}_{[c],[d]})$, Exp. 1
(e) $\mu(U^{s}_{[c],[d]})$, Exp. 2
(f) $\mu(U^{s}_{[c],[d]})$, Exp. 2*
Figure 8: Distributions of the values in the data sets $\mu(\Lambda^{s}_{[c],[d]})$ with $\Lambda\in\{\Theta,U\}$ with breakdown by human-machine interaction dynamics (function $a$), influencing augmenting effects
Figure 9: Experiment 1, mean of the distribution $\mu(U^{s}_{c,d})$, with breakdown by human-machine interaction dynamics (function $a$). The bar plots titled Exp. 1#1\# and Exp. 1£ show the mean of the distribution $\mu(U^{s}_{c,d})$ computed on subsets of $U^{s}_{c,d}$ data with $\delta_{HM}\leq 0.2$ and $\delta_{HM}\geq 0.8$, respectively. In short, Exp. 1#1\# focuses on low growth of margin of contribution over epochs for human-machine skills, while Exp. 1£ focuses on high growth.
(a) Mean of the distributions $\sigma(U^{s}_{c,d})$
(b) Mean of the distributions $\rho(U^{s}_{c,d})$
Figure 10: Mean of the distributions $\omega(U^{s}_{c,d})$ with $\omega\in\{\sigma,\rho\}$, with breakdown by experiment and by human-machine interaction dynamics (function $a$)
(a) Standard deviation of the distributions $\mu(U^{s}_{c,d})$
(b) Range of the distributions $\mu(U^{s}_{c,d})$
Figure 11: Standard deviation and range of the distributions $\mu(U^{s}_{c,d})$, with breakdown by experiment and by human-machine interaction dynamics (function $a$)
@@
Figure 12: Histograms of the distributions of the margins of error for $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$ and $\mu(\Lambda^{s}_{c,d})$ with $\Lambda\in\{\Theta,U\}$
(d) $\Lambda^{s}$
(e) $\Lambda^{s}_{c}$
(f) $\Lambda^{s}_{c,d}$
Figure 13: Scatter plots of $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{[c],[d]})\big)$ versus $\mu(\Lambda^{s}_{[c],[d]})$ with $\Lambda\in\{\Theta,U\}$

### v6

@@ 3 Design of experiment
The setup for our Monte Carlo experiments is defined by a combination of experiment-agnostic and experiment-specific configurations. The experiment-agnostic configuration is made of three components: a general setup, an numeric setup, and a categorical setup. The general setup, common to all experiments, specifies $N$, $E$, $K$, $P_{C}$ and $P_{D}$. In our study, we focus on a single firm ($N=1$) walking through ten epochs ($E=10$). With regards to the number of replications, which remarkably impacts the computational effort required by simulations, we set $K=1000$, conditioning the eventual increase of $K$ to the analysis of the margin of error at 95% confidence level at the end of the first round of experiments. At each replication, the firm is randomly assigned a skill policy according to $P_{c}\sim U\{P_{H},P_{HM},P_{M}\}$, and at each epoch the task difficulty is randomly drawn from $P_{d}\sim U\{P_{Low},P_{Med},P_{High}\}$; the choice of uniformly-distributed probabilities is due to realize an equal split of the simulation resources over all the potential cases under analysis.
The numeric setup (two-decimals precision) is obtained by algorithmic selection of the values of $\gamma_{HM}$, $mc_{H}$, $mc_{M}$, $mc_{HM}$, $\delta_{HM}$, $t_{err}$ and $c_{err}$ optimizing the coverage of the parameter space. With regards to the applicable value-ranges for each numeric parameter, we implement plausible calibrations. For $\gamma_{HM}$, we set the range $[1,2]$, limiting the augmenting factor to increase only, up to +100%. For the unitary margins of contribution associated to each skill policy, we set $mc_{H}\in[0.4,0.6]$ and $mc_{M}\in[0.7,0.9]$ to represent a substantial cost advantage for automation. Then, we model the human-machine approach in terms of a dual-skill policy that incurs in a dual-cost structure defined as $mc_{HM}\in[0.1,0.5]$, with the maximum and minimum limits of the range computed as $mc_{HM}^{max}=1-(1-mc_{H}^{max})-(1-mc_{M}^{max})$ and $mc_{HM}^{min}=1-(1-mc_{H}^{min})-(1-mc_{M}^{min})$. For $\delta_{HM}$ we set the range $[0,1]$, thus imposing a limit of +100% to the growth of the margin of contribution of skill policy $HM$ over epochs. Finally, for both $t_{err}$ and $c_{err}$ we set a $[0,1]$ range, to explore the whole space of possibilities. Based on these constrains, the algorithmic selection is operated by combining the results obtained through the Latin Hypercube Sampling method (LHS) (mckay2000comparison) and the Maximin method (johnson1990minimax) (no duplicates found, for a total of 480 unique numeric setups). Before consolidating the algorithmically-optimized values, we investigate the resulting configuration space for $mc_{c}$. To this end, we compute $\Delta mc_{HM}=mc_{HM}-[1-(1-mc_{H})-(1-mc_{M})]$ to represent the gap between the values of $(mc_{H},mc_{HM},mc_{M})$ selected by LHS and Maximin and the ideal assumption that $mc_{HM}$ should account for the cost of both skills. The distribution of $\Delta mc_{HM}$, significantly deviating from normality, shows a 0.00 mean and a 0.14 standard deviation (see Figure 3(a)). By grouping the values of $mc_{H}$ and $mc_{M}$ in ranges of 0.05 granularity and visualizing the means of $\Delta mc_{HM}$ and $mc_{HM}$ by group (see Figures 3(b) and3(c), respectively) we observe that the actual $mc_{HM}$ deployed in simulations has a comparative advantage versus our ideal hypotheses when both $mc_{H}$ and $mc_{M}$ are in the lower portion of their ranges, and a comparative disadvantage when both $mc_{H}$ and $mc_{M}$ are in the upper portion of their ranges. Nevertheless, $mc_{HM}$ is always lower than $mc_{M}$ by construction, and lower than $mc_{H}$ in 93% of the cases resulting from the selection operated by LHS and Maximin . An overview of the values resulting from LHS and Maximin is synthetically illustrated by means of a pairwise representation in Figure 4. To complete the numeric setup, we impose $\delta_{H}=0$ and $\delta_{M}=0$ as a default, representing the case of human and machine skills being already on their plateau of efficiency.
For the categorical setup, we set the options available for the functions $a$ and $g^{\prime}$. With regards to $a$, we implement three options:
@@
where $\theta_{HM}$, $\theta_{H}$ and $\theta_{M}$ are notational simplifications for $\theta_{i,c=HM,d,e}$, $\theta_{i,c=H,d,e}$ and $\theta_{i,c=M,d,e}$, $\forall(i,d,e)$. Herein, meanmean represents the case of simply averaging the abilities of human and machine skills with no augmenting effects, while collaboratecollaborate and superpowersuperpower represent moderate and strong augmentation, respectively (for notational convenience, these options may also be noted as $HM$, $HM+$ and $HM++$). For the function $g^{\prime}$, we deploy two options:
@@
with $k=5$, where $y$ is a notational simplification for $y_{i,c,d,e}$ and $\theta$ stands for $\theta_{i,c,d,e}$. These options reflect two alternative dynamics for the performance-to-output map.
Finally, the experiment-specific setup involves the hypotheses regarding the distributions $P(\theta)_{c,d,e}$. By means of an overall design choice, we model $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$ to benefit from the range constrained to $[0,1]$ by construction (a natural fit for normalized performance $\theta$) and from its flexibility, allowing us to easily model a variety of skill profiles. With regards to the choice of $\alpha_{c,d,e}$ and $\beta_{c,d,e}$, in the absence of micro-data on skill performance we calibrate to align with stylized conditions. In the first experiment (Exp. 1), we model skill performance such that machines have a comparative advantage in low-complexity tasks, while humans have a comparative advantage in high-complexity task, configuring the human skills to improve across all task difficulties over epochs, while machine skills remain static (training at time $t$ only, with no infra-epoch training) -see Figure 5(a). In a second experiment (Exp. 2), we model a higher performance profile for machine skills with respect to Exp. 1 over medium-to-high-difficulty tasks, as illustrated in Figure 5(b). Finally, in a third experimental session (Exp. 3), we replicate the setup of Exp. 2 while imposing a static human performance over epochs. A comprehensive view of the setup parameters $(\alpha,\beta)_{c,d,e=0}$ and $(\alpha,\beta)_{c,d,e=E-1}$ for all the experiments is detailed in Table 1. In short, the setup of Experiment 1 is calibrated to mimic a state of technological advancement in line with the SOTA at the time this manuscript is written, combined with the idea that human skills may have an evolutionary infra-epoch advantage with respect to machine models. Then, with the Experiment 2 we enhance the technology, making artificial skill outperforming human skills whatever the generalization difficulty involved. Finally, with Experiment 3, we penalize human skills by disabling their ability to evolve over epochs. All the experiments are executed on computing instances deploying Python 3.11, generating a total of 86.4 million in-silico task executions.
(a) Distribution of $\Delta mc_{HM}$
(b) Heatmap of the mean of $\Delta mc_{HM}$
(c) Heatmap of the mean of $mc_{HM}$
Figure 3: Analysis of the actual values determined by LHS and Maximin for $mc_{HM}\in[0.1,0.5]$. In 3(c), each heatmap quadrant reports the ideal range for $mc_{HM}$ in the form [min,max] and the mean of the actual values $mc_{HM}$ determined by LHS and Maximin.
@@
Figure 5: Experimental setup for the distributions of probabilities $P(\theta)_{c,d,e}$ for all skill policies and difficulty levels at epoch $e=0$ and e=E1e=E-1 (Exp. 1 and 2), representing human skills to evolve through epochs, while machine skills are static. A different mix of the same configurations constitutes the basis for Exp. 3.
Table 1: Parameters $(\alpha,\beta)_{c,d,e}$ for the distributions $P(\theta)_{c,d,e}$
Exp. 1 Exp. 2 Exp. 3
$e=0$ e=E1e=E-1 $e=0$ e=E1e=E-1 $e=0$ e=E1e=E-1
$\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$ $\alpha$ $\beta$
$P(\theta)_{c=M,d,e}$ High 2 3 3 2.5 2 3 3 2.5 2 3 2 3
Med 5 5 5 3 5 5 5 3 5 5 5 5
Low 3 2 6 2 3 2 6 2 3 2 3 2
$P(\theta)_{c=H,d,e}$ High 2 4 2 4 6 4 6 4 6 4 6 4
Med 4 2 4 2 8 2.5 8 2.5 8 2.5 8 2.5
Low 10 2 10 2 10 2 10 2 10 2 10 2
@@ 4 Results
Contrary to the conventional wisdom that a human-in-the-loop policy is a safe risk-mitigating strategy, our simulations suggest it prone to contextual economic volatility. To quantitatively characterize this effect, we complement the presentation of the plain results of our individual experiments, denoted as Exp. 1, 2, and 3 (see Figure 6(a)), with close-up perspectives focused on situations characterized by high cost of errors ($t_{err}\leq 0.1$ and $c_{err}\geq 0.8$), denoted as Exp. 1*, 2* and 3* (see Figure 6(b)) and by the low or high ability of the organization to walk the experience curve with the human-machine skill policy over epochs ($\delta_{HM}\leq 0.2$ or $\delta_{HM}\geq 0.8$, respectively), denoted as Exp. 1# and 1£ (see Figure 6(c)).
As shown in Figure 6(a), when confronting high-difficulty tasks in the plausible setting of Experiment 1, the human-machine skill policy achieving augmenting effects reveals higher bottom-line utility than both the human-exclusive and the machine-exclusive strategies. However, when failing to produce augmenting effects, the human-machine strategy results the least economically effective option, derailing into negative utility and potentially destroying value instead of building it. This is far from the result one may expect from a reassuring risk-mitigating approach. Interestingly, the synergic (augmenting) human-machine skill policy is the preferable option in high-difficulty tasks across all the experiments when i) the cost of errors is high (see Figure 6(b)) and ii) regardless of the increase rate over epochs of the margin of contribution associated to the human-machine policy (see Figure 6(c)). This suggests that, in situations where poor task execution triggers expensive shocks, combining human and machine skills can be pivotal even when machine skills become extremely powerful. Also, the increase of the margin of contribution associated to the human-machine skill policy over epochs mitigates the erosion of value, while not safely resolving it (see Figure 6(c), with the human-machine skill policy failing to achieve augmentation closing just at break-even). This highlights the importance of augmentation towards the bottom-line economic effectiveness of the human-machine skill policy, even when machine skills do a great job in high-difficulty tasks.
Next, while denoting the emergent critical finding of the all-or-nothing nature of the human-machine skill policy, our simulation data replicate two relevant and established stylized facts from the existing literature, as presented in Section 1. First, all the experimental results illustrated in Figure 6 (central tendencies) align with the notion that automation tends to be the most economically-effective strategy when task difficulty is low-to-medium, underscoring a superiority in repetitive or compartmentalized tasks. On the other hand, the human-exclusive skill policy can be competitive with respect to automation as long as the human skills can better handling generalization difficulty in situation characterized by high cost of errors. These results are not a straightforward consequence of our design assumptions, because hypotheses are formulated on performance and costs/margins, while our considerations pertain to utility as determined through simulation over the chosen parameter space. Equally interesting is that, while performance is a variable that can be controlled by the decision-maker through careful skill selection or programming, the cost of errors is an exogenous factor.
With regards to dispersion, the mean of the distributions $\sigma(U^{s}_{c,d})$ and $\rho(U^{s}_{c,d})$ presented in Figures 7 and 8 highlights a consistently-lower variability for the values in $U^{s}_{c,d}$ over the simulation space for the human-machine skill policy versus both the human-exclusive or machine-exclusive strategies. In contrast, in Figures 9 and 10 we observe that the standard deviation and the range of the distribution $\mu(U^{s}_{c,d})$ for the human-machine skill policy tends to be often slightly higher than than the one realized by the human or machine skills alone. In both the observations, augmenting effects seem not to have a visible impact. Then, our interpretation is that the hybrid skill strategy contributes more to reducing the variability of utility over the spectrum of scenarios under consideration, while slightly but less evidently increasing the variability of $\mu(U^{s}_{c,d})$. This may explain why the combination of human and machine skills can be perceived as a risk-mitigating strategy, because the reduction of variability within a specific context naturally suggests lower risks; however, the variability alone is not sufficient to avoid the risk of being near or below utility break even when the variables describing the context are subject to uncertainty.
A more fine-grained detail of our results is enclosed in the Appendix of this manuscript. In Table 4, 5 and 6 we present the description of the distributions $\omega(\Lambda^{s}_{c,d})$ with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,U\}$ through their arithmetic mean, standard deviation, range (max-min), and median value. In Figures 14 and 15 we provide an intuitive visualization of the distributions $\mu(U^{s}_{c,d})$ and $\sigma(U^{s}_{c,d})$ by means of overlapping violin and box plots, with breakdown by skill/interaction policy and task difficulty for each experiment and focused perspective.
Figure 6: Mean of the distributions $\mu(U^{s}_{c,d})$ by experiment and by skill/interaction policy
Figure 7: Mean of the distributions $\sigma(U^{s}_{c,d})$ by experiment and by skill/interaction policy
Figure 8: Mean of the distributions $\rho(U^{s}_{c,d})$ by experiment and by skill/interaction policy
Figure 9: Standard deviation of the distributions $\mu(U^{s}_{c,d})$ by experiment and by skill/interaction policy
Figure 10: Range of the distributions $\mu(U^{s}_{c,d})$ by experiment and by skill/interaction policy
@@ 5 Discussion
To allow the reader a more precise understanding of the margins of error of $\mu(\Theta^{s}_{[c],[d]})$ and $\mu(U^{s}_{[c],[d]})$ obtained with $K=1000$ as estimates of the true mean ($K\rightarrow\infty$), for each simulation $s$ of all the experiments we compute the estimated MOE at 95% confidence level for $\mu(\Lambda^{s})$, $\mu(\Lambda^{s}_{c})$ and $\mu(\Lambda^{s}_{c,d})$ with $\Lambda\in\{\Theta,U\}$. Formally, $\forall s$, we compute $\text{MOE}_{95}\big(\mu(\Lambda^{s})\big)=1.96*\sigma(\Lambda^{s})/\sqrt{K^{s}}$, $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c})\big)=1.96*\sigma(\Lambda^{s}_{c})\sqrt{K^{s}_{c}}$, and $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{c,d})\big)=1.96*\sigma(\Lambda^{s}_{c,d})\sqrt{K^{s}_{c,d}}$ with $K^{s}=K*E=10000$, and $K^{s}_{c}$ and $K^{s}_{c,d}$ globally and prudentially set to $3000<K*E/|C|\sim 3333$ and $1000<K*E/(|C|*|D|)\sim 1111$, respectively. We acknowledge that, within each simulation $s$, the sets $\Lambda^{s}$, $\Lambda^{s}_{c}$ and $\Lambda^{s}_{c,d}$ pool values that are related to different epochs $e$, thus potentially in relationship with different Beta distributions for skill policies $H$ and $HM$ (this is due to the deterministic interpolation of the parameters $\alpha$ and $\beta$ over epochs for the skill $H$), meaning that the data is heterogeneous and consists of independent samples of smaller size. However, the margins of error computed on the pooled standard deviation represent a conservative upper bound for the granular error that should be selectively computed for each epoch, because the pooled standard deviation combines the actual sampling uncertainty with the additional variability introduced by the parameter interpolation. The distributions of $\text{MOE}_{95}\big(\mu(\Lambda^{s}_{[c],[d]})\big)$ with $\Lambda\in\{\Theta,U\}$ are described in Figure 11 and in Table 2. with mean consistently below 0.01.
@@
Figure 11: Distributions of the margins of error $\text{MOE}_{95}(\Lambda^{s}_{[c],[d]})$ with $\Lambda\in\{\Theta,U\}$